5. Complete the following Balance Sheet based on the given data Total Debt to Owners Equity is 0.6 Current Debt to Total Debt is 0.4 Fixed Asset to Owners Equity is 0.6 Total Asset Turnover is 2 Inventory Turnover is 8 Gross profit 40% .

Cash............

Inventory........

Total current asset....

Fixed asset...........

Current liability...........

Long term debt..........

Total debt..............

Owner's equity..........1000

Total liability and owner's equity...


1
Expert's answer
2022-01-17T12:42:15-0500

Total Debt to Owners Equity=0.6

Let Total debt= X

"\frac{X}{Owners Equity}=0.6"

"\frac{X}{1000}=0.6"

X= 600

Total Debt =600

Current Debt to Total Debt=0.4

"\frac{Current Debt}{Total Debt}=0.4"

Current Debt="0.4\times 600= 240"


Fixed Assets to Owners Equity

"\frac{Fixed Assets}{Owners Equity}=0.6"

FA= 0.6"\times 1000= 600"


Long Term Debt= Total debt - Current Debt

= 600-240= 360

"Debt ratio=\frac{Total Debts}{Total Assets}"


0.4="\frac{600}{Total Assets}"

Total Assets= "\frac{600}{0.4}= 1500"


Current Assets= Total Assets- Fixed Assets

= 1500- 600= 900

Total Assets turn over="\frac{Sales}{Total Assets}"

Sales="1500\times2= 3000"

Gross Profit margin="\frac {Gross Profit}{Sales}"

Sales= 40%"\times" 3000= 1200

Cost of goods sold= Sales - Gross Profit

= 3000-1200= 1800

Inventory Turnover= "\frac{Cost of goods Sold}{Average Stock}"

Average Stock="\frac{Cost of goods Sold}{Inventory turnover}"


"\frac{1800}{2}= 900"

Total Stock= "\frac{900}{2}= 450"

Cash= Current Assets- Stock

= 900-450= 450